OpenStudy (anonymous):

howyo solve this y=2 over5 x+2

OpenStudy (anonymous):

y=2/5x+2

OpenStudy (tcp2315):

you need to graph this

OpenStudy (anonymous):

yes

OpenStudy (tcp2315):

so are you asking how to graph it?

OpenStudy (anonymous):

if you could help me with the points i can get it from there

OpenStudy (tcp2315):

sorry for the long answer, but just find the y intercept and the x intercept. you do this by setting x to be zero (for the yint) and setting y to be zero for the x intercept. x int: 0=2/5x+2 -2=2/5x -5=x So the x intercept is the point (-5,0) y int: y=2/5(0) +2 y=2 so the y intercept is (0, 2) There you go, two points!